8. A bag contains 4 red marbles, 3 yellow marbles, 2 black marbles and 1 green marbles. You select a marble at random. Find each probability.
a) Red
b) Yellow
c) Black
d) Green

Answer by Mathtut(3670)

A bag contains 4 red marbles, 3 yellow marbles, 2 black marbles and 1 green marbles. You select a marble at random. Find each probability.
a) Red
b) Yellow
c) Black
d) Green
a)red: 4/10=2/5
b)yellow:3/10
c)Black:2/10=1/5
d)Green:1/10